# Understanding Impact of Human Feedback via Influence Functions

### v1

@@ 5.1.2 Results and analysis
The ROC curves in Figure 2 demonstrate that our method, utilizing influence functions, significantly outperforms all baselines in detecting length and sycophancy biases. It achieves AUC values of 0.8 for length bias and 0.711 for sycophancy bias, compared to 0.6 for other threshold-based detectors. Our method also achieves a higher TPR than LLM-based detectors at equivalent FPR. Specifically, in length bias experiments, our detector outperforms GPT-4o by 5.3% and Gemini-1.5-Pro by 25.6%. For sycophancy bias, it exceeds GPT-4o by 14.8% and Gemini-1.5-Pro by 11.9%. On average, our method identifies 14.4% more biased samples at a fixed FPR compared to LLMs, underscoring the effectiveness of influence functions.
Furthermore, we note that length bias is easier to detect than sycophancy bias across all methods. Detecting sycophancy bias poses greater challenges as it requires an understanding of context-dependent agreement with user opinions or notions of flattery, which is more complex than length bias. Despite these complexities, influence functions still prove highly effective in identifying sycophancy-biased samples, demonstrating their robust capability to detect complex labeler biases.
@@
In Figure 3, we present a qualitative analysis of the most positively-contributing and negatively-contributing samples for both length and sycophancy bias experiments. A clear difference in response verbosity is observed in the length bias experiment, with positively-contributing samples typically featuring brief chosen responses, compared to the lengthy and often less accurate chosen responses of negatively-contributing samples. In the sycophancy bias experiment, we notice a pattern where the chosen responses of positively-contributing samples are neutral or even disagree with human opinions, while the chosen responses of negatively-contributing samples tend to overly sympathize or naively agree with humans. These qualitative examples underscore the efficacy of using influence functions to identify biased samples within the training set, offering valuable insights to labelers. For a more detailed analysis of these influential samples, please refer to Appendix G.
Figure 3: Most positively-contributing (left) and negatively-contributing (right) responses from the length bias experiment (top) and the sycophancy bias experiment (bottom). In the length bias experiment, positively-contributing chosen responses are concise, while negatively-contributing chosen responses are longer and often off-topic. In the sycophancy bias experiment, positively-contributing chosen responses are objective and helpful, whereas negatively-contributing chosen responses are excessively sympathetic.
Figure 5: Runtime comparison for different training dataset sizes. Our method is 2.5 times faster, requiring 11.1 hours compared to DataInf’s 28.8 hours for a $10^{5}$ sized training dataset.
@@
We also investigate the impact of validation set size for influence functions and the number of few-shot examples for LLM baselines in Appendix H. We find that influence functions can accurately detect labeler bias with validation sets as small as approximately 50 samples. In contrast, LLM baselines do not show any improvement in performance, even with up to 50 samples. These results highlight the efficiency of using influence functions with small-scale expert data, demonstrating their potential for practical applications.
@@
To verify the computational efficiency of our method, we compare the runtime of our approach to DataInf (Kwon et al., 2024) across various training dataset sizes while using reward models of the same size and keeping the validation set size fixed at 1,000 samples. Figure 5 shows that our method is approximately 2.5 times faster than DataInf. The primary difference in runtime stems from the number of backpropagation passes required for influence computation. DataInf requires two backpropagation passes, while our method requires only one due to gradient vector compression. [^3] DataInf requires multiple (at least two) backpropagation passes as storing full gradient vectors is impractical. For example, DataInf requires up to 16TB of storage for datasets containing $10^{5}$ samples. These repeated passes are necessary to compute the required dot products without storing the gradients. While compression in our method takes 11.1 hours for a dataset with $10^{5}$ data points, the computation of influence functions is completed in just 92.3 seconds after compression. In contrast, DataInf, which does not apply compression, requires two backpropagation passes and cannot store gradient vectors efficiently, resulting in a runtime of 28.8 hours for the same dataset.
5.2 Labeling strategy oversight using influence functions
@@ 5.2.1 Experimental setup
We provide an overview of our labeler strategy oversight experiment in Figure 6, which illustrates a scenario designed to model simulated labelers and their labeling strategies. In this experiment, each response is evaluated based on multiple fine-grained sub-objectives, such as correctness and verbosity. Labelers evaluate the overall score of a response using a weighted sum of sub-objectives, formulated as $r=\textbf{w}^{\top}(r_{1},r_{2},r_{3},r_{4})$, where each $r_{i}\in\mathbb{R}$ represents a sub-objective score of a response. We assume that the sub-objective scores are consistent across labelers, but the weight vector $\textbf{w}\in\mathbb{R}^{4}$, which represents a labeler’s strategy for prioritizing these sub-objectives, varies among them. To generate feedback, labelers determine the preference label $z$ by comparing the scores of two responses, $z=\mathbb{I}(\textbf{w}^{\top}\textbf{r}^{(0)}<\textbf{w}^{\top}\textbf{r}^{(1)})$, where $\textbf{r}^{(0)}$ and $\textbf{r}^{(1)}$ are the sub-objective score vectors for the response pairs $y^{(0)}$ and $y^{(1)}$. This framework enables us to simulate different labeler strategies effectively.
@@
Figure 6: An overview of our labeling strategy oversight experiment. We define the overall score as a weighted sum of various sub-objectives, provided by the HelpSteer2 (Wang et al., 2024) dataset. Alice and Bob labels binary preference $z_{\tt A},z_{\tt B}$ between responses using their respective labeling strategy $\textbf{w}_{\tt A},\textbf{w}_{\tt B}$ Influence functions are estimated upon Alice’s validation set $\mathcal{D}_{\tt A}$, identifying redundant or potentially detrimental samples in $\mathcal{D}_{\tt B}$. This information is used to update Bob’s labeling strategy $\textbf{w}_{\tt B}$ by applying a support vector machine.
@@
To update Bob’s labeling strategy, we first identify samples that most positively and negatively impact his labeling accuracy compared to Alice, using influence functions. Given a learned reward model $r_{\theta}$, the influence value $\mathcal{I}_{\tt val}(\mathbf{d}_{i})$ is calculated for each data point $\mathbf{d}_{i}\in\mathcal{D}_{\tt B}$ based on $\mathcal{L}_{\tt val}(\mathcal{D}_{\tt A};\theta)$. Samples with an influence score $\mathcal{I}_{\tt val}(\mathbf{d}_{i})$ exceeding a specified threshold are classified as negatively contributing, while those below the threshold are deemed positively contributing. We then update weights by classifying these positive and negative samples based on their sub-objective scores using support vector machines (Cortes & Vapnik, 1995). Details on the weight updates are provided in Appendix F. Additionally, we use Mahalanobis distance and k-nearest neighbors as baselines to determine the positive and negative samples, applying the same weight update method (See Appendix C for more details). [^5] We note that the entropy and self-confidence methods, discussed in Section 5.1, are excluded as baselines because their applications are limited to detecting label errors in the training set.
@@
We evaluate the performance of weight updates (i.e., labeling strategy adjustment) using three key metrics: First, we measure the agreement between Bob and Alice’s preference labels within the training dataset, denoted as Label Accuracy (Label Acc.). Additionally, we report the validation accuracy of the reward model trained on $\mathcal{D}_{\tt B}$, referred to as Reward Model Accuracy (RM Acc.). Finally, we calculate the cosine similarity between $\textbf{w}_{\tt A}$ and $\textbf{w}_{\tt B}$ to assess how closely Bob’s strategy aligns with Alice’s expert strategy, noted as Cosine Similarity (Cos Sim.).
@@ 5.2.2 Results and analysis
To further examine the impact of using a small validation set, we present performance metrics across different validation set sizes starting from 10 samples. Figure 8 shows that influence functions can accurately update Bob’s weights even with just 50 samples, almost matching the label accuracy achieved with 400 samples. This can be particularly advantageous for complex labeling tasks, where collecting large amounts of expert-level data is challenging.
Limitations
We highlight several constraints in our experimental setup that may not extend to real-world settings. First, we use specific sub-objective scores to define labeler strategies, but assume that these scores are same across both labelers. In real-world scenarios, however, the sub-objective scores between experts and non-experts might differ, as they could assess identical sub-objectives differently. Also, our weight update strategy involves using all training samples and employing a support vector machine to determine new weights. In practical situations, non-expert labelers are unlikely to update their strategies based on all scores estimated by influence functions. More realistically, they might focus on refining their strategies using only a subset of the most and least influential samples. Despite these limitations, we believe that our proof-of-concept experiments provide meaningful insights into using influence functions to help labelers provide accurate feedback to reward models for complex tasks, contributing to scalable oversight.
@@ 6 Conclusion
While this research does not explicitly showcase examples from preference datasets containing offensive or harmful content, we want to notify readers of the possibility that such instances may exist in datasets we release through supplementary materials. During manual inspection of samples from the helpful split of Anthropic’s Helpfulness-Harmlessness dataset (Bai et al., 2022a), we observed a few examples containing swear words, though they were limited in number. Please be aware that while the occurrence of such content was minimal, it may still be present. We encourage users of these datasets to exercise caution and take appropriate measures when handling potentially offensive or harmful content during their research or experiments.
Reproducibility statement
In order to facilitate the reproducibility of our work, we provide our code with detailed instructions to ensure that all key elements of our experiments can be replicated. Specifically, we provide supplementary materials that include anonymous links to datasets that are used in our experiments for Section 5.1 and Section 5.2. The code for our experiments is provided in the supplementary files. These resources ensure that our results can be reproduced.
@@ A.1 Vector compression method
Dataset Size Storage Requirement Before Compression (GB) Storage Requirement After Compression (GB)
@@
In this section, we describe the vector compression method employed in our work: the one-permutation, one-random-projection (OPORP) technique (Li & Li, 2023). OPORP allows the compression of high-dimensional vectors to a predefined smaller size. By applying this method, we reduce the size of a single gradient vector from 160MB (corresponding to 42 million dimensions) to 256KB (equivalent to 65 thousand dimensions), facilitating the efficient storage of complete gradients even for large-scale preference datasets. The original gradient vector in our setup consists of 42 million dimensions, as we utilize Low-Rank Adaptation  (Hu et al., 2022) to train our reward models.
@@ A.2 Performance comparison with datainf
Appendix B Datasets detail
@@ B.1 Bias detection
Experiment Dataset Source Corruption ratio Size (Train) Size (Validation)
Length bias Anthropic-HH (helpful) 6.56% 15000 6121
Sycophancy bias Anthropic-HH (helpful-online) 4.17% 15000 1071
Table 3: Details on datasets used in bias detection
We use Anthropic’s Helpfulness-Harmlessness dataset(Anthropic-HH(Bai et al., 2022a) for bias detection experiments. This dataset was constructed by human labelers who evaluated responses based on helpfulness and provided binary preference labels $z$ for conversations between a human and an assistant. Table 3 summarizes dataset information in this experiment.
@@
We randomly sampled 15k samples from Anthropic-HH-helpful dataset, the helpful split of Anthropic-HH dataset, where responses were evaluated regarding helpfulness. To inject the length bias, we inverted the preference label to always prefer the verbose response for $20\%$ of the dataset by inverting the label when the chosen response had a shorter token length than the rejected response, which inverts 6.56% of the dataset. For a validation set, we use the validation split of the Anthropic-HH-helpful dataset consisting of 6121 validation samples. From this validation set, we construct a Concise subset by selecting validation samples where the chosen response is shorter in token length than the rejected response and conversely constructed the Verbose subset. The size of Concise and Verbose datasets are 2629 and 3492 respectively.

### v2

@@ 5.1.2 Results and Analysis
The ROC curves in Figure 2 demonstrate that our method, utilizing influence functions, significantly outperforms all baselines in detecting length and sycophancy biases. It achieves AUC values of 0.8 for length bias and 0.711 for sycophancy bias, compared to 0.6 for other threshold-based detectors. Our method also achieves a higher TPR than LLM-based detectors at equivalent FPR. Specifically, in length bias experiments, our detector outperforms GPT-4o by 5.3% and Gemini-1.5-Pro by 25.6%. For sycophancy bias, it exceeds GPT-4o by 14.8% and Gemini-1.5-Pro by 11.9%. These results underscore the effectiveness of using influence functions for bias detection.
We also observe that length bias is easier to detect than sycophancy bias across all methods. Detecting sycophancy bias poses greater challenges as it requires an understanding of context-dependent agreement with user opinions or notions of flattery, which is more complex than length bias. Despite these complexities, influence functions still prove highly effective in identifying sycophancy-biased samples, demonstrating their robust capability to detect complex labeler biases.
Additionally, we assess the validation accuracy of the retrained reward model after modifying the training data by flipping the preference labels of negatively-contributing samples. Even with a small number of label flips, validation accuracy improves significantly on the Concise and Less Sycophantic sets (see Table 10 and Appendix G for details).
@@
In Appendix H, we present a qualitative analysis of the most positively-contributing and negatively-contributing samples for both length and sycophancy bias experiments. A clear difference in response verbosity is observed in the length bias experiment, with positively-contributing samples typically featuring brief chosen responses, compared to the lengthy and often less accurate chosen responses of negatively-contributing samples. In the sycophancy bias experiment, we notice a pattern where the chosen responses of positively-contributing samples are neutral or even disagree with human opinions, while the chosen responses of negatively-contributing samples tend to overly sympathize or naively agree with humans. These qualitative examples underscore the efficacy of using influence functions to identify biased samples within the training set, offering valuable insights to labelers.
@@
We conduct ablation studies to analyze the impact of validation set composition, size and distribution in  Appendix I. In Figure 11, we find that our Concise and Verbose set performs optimally compared to their counterparts (Verbose and More Sycophantic, see Appendix I for details), as well as the Full validation set. These results highlight the importance of constructing targeted validation sets, as noted in Remark 4.1. Additionally, Figure 12 show that influence functions can accurately detect labeler bias with validation sets as small as 50 samples. In contrast, as shown in Figure 13, LLM baselines exhibit no improvement in performance, even when provided with up to 50 samples. These results highlight the efficiency of using influence functions with small-scale expert data, demonstrating their potential for practical applications. In Table 11, we further investigate the impact of distributional differences between training and validation sets, showing that using a validation set from HelpSteer2 Wang et al. (2024) yields a suboptimal AUC of 0.620. This underscores the necessity of closely aligning validation and training distributions to achieve optimal performance.
Figure 3: Runtime comparison for different training dataset sizes. Our method is 2.5 times faster compared to DataInf.
@@
To verify the computational efficiency of our method, we compare the runtime of our approach to DataInf (Kwon et al., 2024) across various training dataset sizes while using reward models of the same size and keeping the validation set size fixed at 1,000 samples. Figure 3 shows that our method is $\sim$2.5 times faster than DataInf. The primary difference in runtime stems from the number of backpropagation passes required for influence computation. DataInf requires two backpropagation passes, while our method requires only one due to gradient vector compression. [^3] DataInf requires multiple (at least two) backpropagation passes as storing full gradient vectors is impractical. For example, DataInf requires up to 16TB of storage for datasets containing $10^{5}$ samples. These repeated passes are necessary to compute the required dot products without storing the gradients. While compression in our method takes 11.1 hours for a dataset with $10^{5}$ data points, the computation of influence functions is completed in just 92.3 seconds after compression. In contrast, DataInf, which does not apply compression, requires two backpropagation passes and has a runtime of 28.8 hours.
Detecting Biased Samples in the Original Human Dataset
While our primary analysis uses a dataset with flipped labels to simulate labeler bias, we extend our method to the original Anthropic-HH dataset to evaluate its effectiveness on real data. Using a human survey, we examine the preference labels of the top 100 negatively-contributing samples (Top-100), selected based on influence function scores from the Concise set, which targets length biased samples. For comparison, we also inspect a randomly selected set of 100 samples (Random-100). As shown in Table 1, 47% of samples in the Top-100 set are mislabeled, compared to 13% in the Random-100 set. This substantial gap demonstrates that our method effectively identifies mislabeled samples in real datasets by detecting labeler bias. Moreover, it highlights that labeler biases are not uncommon, reinforcing the practical utility of our approach. For annotation details, results on the sycophancy validation set, and examples of Top-100 samples, we refer readers to Appendix J.
Subset Mislabeled Correct Tie
Top-100 47 38 15
Random-100 13 69 18
Table 1: Human survey results comparing the Top-100 subsets selected using our influence function approach with Random-100 subsets. Significant portions of Top-100 samples are mislabeled compared to Random-100.
5.2 Labeling Strategy Oversight
@@ 5.2.1 Experimental Setup
We provide an overview of our labeler strategy oversight experiment in Figure 4, which illustrates a scenario designed to model simulated labelers and their labeling strategies. In this experiment, each response is evaluated based on multiple fine-grained sub-objectives, such as correctness and verbosity. Labelers evaluate the overall score of a response using a weighted sum of sub-objectives, formulated as $r=\textbf{w}^{\top}(r_{1},r_{2},r_{3},r_{4})$, where each $r_{i}\in\mathbb{R}$ represents a sub-objective score of a response. We assume that the sub-objective scores are consistent across labelers, but the weight vector $\textbf{w}\in\mathbb{R}^{4}$, which represents a labeler’s strategy for prioritizing these sub-objectives, varies among them. To generate feedback, labelers determine the preference label $z$ by comparing the scores of two responses, $z=\mathbb{I}(\textbf{w}^{\top}\textbf{r}^{(0)}<\textbf{w}^{\top}\textbf{r}^{(1)})$, where $\textbf{r}^{(0)}$ and $\textbf{r}^{(1)}$ are the sub-objective score vectors for the response pairs $y^{(0)}$ and $y^{(1)}$. This framework enables us to simulate different labeler strategies.
@@
Figure 4: An overview of our labeling strategy oversight experiment. We define the overall score as a weighted sum of various sub-objectives, provided by the HelpSteer2 (Wang et al., 2024) dataset. Alice and Bob label binary preference $z_{\tt A},z_{\tt B}$ between responses using their respective labeling strategies $\textbf{w}_{\tt A},\textbf{w}_{\tt B}$. Influence functions are estimated upon Alice’s validation set $\mathcal{D}_{\tt A}$, identifying redundant or potentially detrimental samples in $\mathcal{D}_{\tt B}$. This information is used to update Bob’s labeling strategy $\textbf{w}_{\tt B}$ by applying a support vector machine.
@@
To update Bob’s labeling strategy, we first identify samples that most positively and negatively impact his labeling accuracy compared to Alice, using influence functions. Given a learned reward model $r_{\theta}$, the influence value $\mathcal{I}_{\tt val}(\mathbf{d}_{i})$ is calculated for each data point $\mathbf{d}_{i}\in\mathcal{D}_{\tt B}$ based on $\mathcal{L}_{\tt val}(\mathcal{D}_{\tt A};\theta)$. Samples with an influence score $\mathcal{I}_{\tt val}(\mathbf{d}_{i})$ exceeding a specified threshold are classified as negatively contributing, while those below the threshold are deemed positively contributing. We then update weights by classifying these positive and negative samples based on their sub-objective scores using support vector machines (Cortes and Vapnik, 1995). Details on the weight updates are provided in Appendix F. Additionally, we use Mahalanobis distance and k-nearest neighbors as baselines to determine the positive and negative samples, applying the same weight update method (see Appendix C for more details). [^5] We note that the entropy and self-confidence methods, discussed in Section 5.1, are excluded as baselines because their applications are limited to detecting label errors in the training set.
@@
We evaluate the performance of weight updates (i.e., labeling strategy adjustment) using three key metrics. Label Accuracy (Label Acc.) measures the agreement between Bob and Alice’s preference labels in the training dataset. Reward Model Accuracy (RM Acc.) refers to the validation accuracy of the reward model trained on $\mathcal{D}_{\tt B}$. Cosine Similarity (Cos Sim.) quantifies the alignment between Bob’s strategy $\textbf{w}_{\tt B}$ and Alice’s expert strategy $\textbf{w}_{\tt A}$.
@@ 5.2.2 Results and Analysis
To further examine the impact of using a small validation set, we report performance metrics across different validation set sizes starting from 10 samples. Figure 14 shows that influence functions can accurately update Bob’s weights even with just 50 samples, implying that our method can be advantageous for complex labeling tasks where expert-level data is scarce.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank Juyong Lee, Kyuyoung Kim, and Roger Grosse for helpful comments on the project. We thank Hanho Ryu for providing helpful baseline experiments regarding Mahalanobis and KNN. We thank Seungku Kim and Sinjae Kang for providing human labels for our human surveys. This work was supported by National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259), Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)). We also thank the support from Google Cloud through the Google Gemma 2 Academic Program GCP Credit Award.
Limitations
Estimation Error
Our approach is based on DataInf which approximates the inverse hessian in influence functions (Kwon et al., 2024). The total approximation error of DataInf is upper-bounded by $\sum_{l=1}^{L}O(d_{l}^{2})$, where $d_{l}$ represents the parameter size of layer $l$, suggesting that influence estimation error may increase as model size increases. However, this bound represents only a worst-case scenario, and the actual error can only determined through comparisons with exact influence estimation methods, an approach that is computationally infeasible for large language models. Instead, we empirically validate our method by evaluating its effectiveness in practical applications such as bias detection and labeler strategy oversight. Our experiments on Llama-3-8B demonstrate that our approach performs well in these tasks, though estimation error may increase for larger models with significantly more parameters.
Bias Detection
We mainly focus on the Anthropic-HH (Bai et al., 2022) dataset, and two prevalent types of bias (length and sycophancy) to verify our approach. While we observe robust performance on different types of bias, additional experiments may be necessary to empirically assess our method’s applicability to different datasets and biases. Our experiments also rely on a fixed validation set requiring manual curation. To reduce this dependency, future research should explore automating validation set curation, building on recent studies showing that frontier AI models can effectively monitor and classify diverse objectives Tan et al. (2024); Baker et al. (2025).
Labeling Strategy Oversight
We highlight several constraints in our experimental setup that may not extend to real-world settings. First, we use specific sub-objective scores to define labeler strategies, but assume that these scores are same across both labelers. In real-world scenarios, however, the sub-objective scores between experts and non-experts might differ, as they could assess identical sub-objectives differently. Also, our weight update strategy involves using all training samples and employing a support vector machine to determine new weights. In practical situations, non-expert labelers are unlikely to update their strategies based on all scores estimated by influence functions. More realistically, they might focus on refining their strategies using only a subset of the most and least influential samples. To address these limitations, future research should focus on scenarios involving multiple labelers, explore non-linear functions to define labeler strategies, and prioritize human-in-the-loop experiments. Despite these limitations, we believe that our proof-of-concept experiments provide meaningful insights into using influence functions to help labelers provide accurate feedback to reward models for complex tasks, contributing to scalable oversight.
@@ Acknowledgments
While this research does not explicitly showcase examples from preference datasets containing offensive or harmful content, we want to notify readers of the possibility that such instances may exist in datasets we release through supplementary materials. During manual inspection of samples from the helpful split of Anthropic’s Helpfulness-Harmlessness dataset (Bai et al., 2022), we observed a few examples containing swear words, though they were limited in number. Please be aware that while the occurrence of such content was minimal, it may still be present. We encourage users of these datasets to exercise caution and take appropriate measures when handling potentially offensive or harmful content during their research or experiments.
@@ Appendix A Vector Compression Details
In this section, we describe the one-permutation, one-random-projection (OPORP) compression method in detail in Section A.1, explaining how it efficiently compresses high-dimensional gradient vectors while preserving dot products. We then compare our approach with DataInf in Section A.2, demonstrating that our method achieves comparable accuracy in influence estimation with significantly improved computational efficiency.
@@ A.1 Vector Compression Method
Data Size Pre-compression (GB) Post-compression (GB)
@@
In this section, we describe the vector compression method employed in our work: the one-permutation, one-random-projection (OPORP) technique (Li and Li, 2023). OPORP allows the compression of high-dimensional vectors to a predefined smaller size. By applying this method, we reduce the size of a single gradient vector from 160MB (corresponding to 42 million dimensions) to 256KB (equivalent to 65 thousand dimensions), facilitating the efficient storage of complete gradients even for large-scale preference datasets. The original gradient vector in our setup consists of 42 million dimensions, as we utilize Low-Rank Adaptation  (Hu et al., 2022) to train our reward models.
@@ A.2 Performance Comparison with Datainf
Appendix B Dataset Details
@@ B.1 Bias Detection
Dataset Characteristics
Experiment Data split Corruption ratio
Length helpful 6.56%
Sycophancy helpful-online 4.17%
Experiment Size (Train) Size (Validation)
Length 15000 6121
Sycophancy 15000 1071
Table 4: Details on datasets used in bias detection. The “Data split” indicates the dataset split of Anthropic-HH.
We use Anthropic’s Helpfulness-Harmlessness dataset (Bai et al., 2022), Anthropic-HH, for bias detection experiments. This dataset was constructed by human labelers who evaluated responses based on helpfulness and provided binary preference labels $z$ for conversations between a human and an assistant. Table 4 summarizes dataset information in this experiment.
@@
We randomly sampled 15k samples from Anthropic-HH-helpful dataset, the helpful split of Anthropic-HH dataset, where responses were evaluated regarding helpfulness. To inject length bias, we inverted the preference label to always prefer the verbose response for $20\%$ of the dataset by inverting the label when the chosen response had a shorter token length than the rejected response, which inverts 6.56% of the dataset. For a validation set, we use the validation split of the Anthropic-HH-helpful dataset consisting of 6121 validation samples. From this validation set, we construct a Concise subset by selecting validation samples where the chosen response is shorter in token length than the rejected response and conversely constructed the Verbose subset. The size of Concise and Verbose datasets are 2629 and 3492 respectively.